Half equations for the titration of ferrocyanide and permanganate in acidic neutral and basic solutions?

acidic: MnO4 ^1- + Fe(CN)6 ^4- =

netural: MnO4 ^1- + Fe(CN)6 ^4- +H2O+ (NH4)2SO4+MnO2 =

basic: MnO4 ^1- + Fe(CN)6 ^4- + NaOH=

for the basic soultion I know one product is MnO4 ^-2 but I'm unsure how to set up their default equations and half equations.
1
Expert's answer
2019-03-08T07:56:22-0500

Solution.

1) acid: MnO4 ^1-+ 5 Fe(CN6) ^4- + 8H ^1+ = 5 Fe(CN6) ^3- + Mn ^2+ + 4H2O

2) netural: MnO4 ^1- + 3 Fe(CN6) ^4- + 2H2O = 3 Fe(CN6) ^3- + MnO2 + 4OH ^1-

3) basic: MnO4 ^1- + Fe(CN6)^4- = MnO4 ^2-+ Fe(CN6) ^3-

Answer:

1) acid: MnO4 ^1-+ 5 Fe(CN6) ^4- + 8H ^1+ = 5 Fe(CN6) ^3- + Mn ^2+ + 4H2O

2) netural: MnO4 ^1- + 3 Fe(CN6) ^4- + 2H2O = 3 Fe(CN6) ^3- + MnO2 + 4OH ^1-

3) basic: MnO4 ^1- + Fe(CN6)^4- = MnO4 ^2-+ Fe(CN6) ^3-

I am sorry, there was error in the design of the answer again.